OpenStudy (camerondoherty):

An equation is shown below: 7x + y = 5 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (anonymous):

A: rearrange the terms, we get \(y = 5 - 7x\). so, for every \(x\in(-\infty,\infty)\), there is a corresponding 'y'. This gives you an infinite pairs of (x,y) that satisfy the given equation.

OpenStudy (anonymous):

B. choose any three values of 'x'. Let us say, -1,0 and 3 \[\begin{array} xx=-1\qquad&y=5-7(-1)=5+7=12\\ x=0&y=5-7(0)=5-0=5\\ x=3&y=5-7(3)=5-21=-16 \end{array}\] so, we have the three points as:\((-1,12),(0,5),(3,-16)\)

OpenStudy (anonymous):

aslso, take a look at this post: http://openstudy.com/study#/updates/52bba173e4b01cdca44a0b0c

OpenStudy (camerondoherty):

Sorry @electrokid i was doing something and thank you!